how do you solve the problem 5m^2=125
Found 3 solutions by jim_thompson5910, Alan3354, Onmyoji.S:
Answer by jim_thompson5910(35256)

5m^2 = 125

5m^2 - 125 = 0

5(m^2 - 25) = 0

5(m - 5)(m + 5) = 0

m - 5 = 0 or m + 5 = 0

m = 5 or m = -5

Answer by Alan3354(69443)

how do you solve the problem 5m^2=125
Divide by 5
m^2 = 25
m = -5, +5

Answer by Onmyoji.S(32)

5mē=125
mē=25
m=-5,m=5